OpenStudy (anonymous):

solve the equation on the interval [0, 2pi): cos x + 2 cos x sin x = 0

jimthompson5910 (jim_thompson5910):

I'll get you started cos(x) + 2*cos(x)*sin(x) = 0 cos(x)( 1 + 2*sin(x) ) = 0 cos(x) = 0 or 1 + 2*sin(x) = 0 You need to solve each of those 2 equations above to get your set of solutions.